A grandmother prepared bowls of fruits for her family. All but 5 are with cherries, all but 4 are with apricots, all but three are with red currants. How many bowls are there in total?
Answer by ptaylor(2198)

Let x=number of bowls in total
x-5=bowls with cherries
x-4=bowls with apricots
x-3=bowls with red currants
(x-5)+(x-4)+(x-3)=x
3x-12=x
2x=12
x=6 bowls in all
CK
6-5=1 bowl with cherries
6-4=2 bowls with apricots
6-3=3 bowls with Currants

Hope this helps---ptaylor